OpenStudy (anonymous):

i want to go with c Which is the equation of the power function p(x) = ax^2 if p(3) = 36? (Points : 1) p(x) = 3x^2 p(x) = 4x^2 p(x) = 9x^2 p(x) = 12x^2

OpenStudy (agent0smith):

p(x) = ax^2 p(3) = 36 so when x=3, p is 36... 36 = a(3)^2 find a

OpenStudy (anonymous):

36 / 3 = a(^2) ? 12 = a^2 px = 12x^2?

OpenStudy (agent0smith):

3^2 is 9.

OpenStudy (anonymous):

4 = a

OpenStudy (anonymous):

the answer is b